APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday August 9, 2004 (MJD 53227) ---=== OBSERVING TEAM ===--- (swing) Jurek Krzesinski (night) Howard Brewington, Steph Snedden ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Declared plates 1403 and 1108 done. And, we got some siginal on two other ones, i.e., 1025, 1144. We had some guider problems, and Landru locked-up. ---=== OBSERVING LOG ===--- Afternoon: ---------- DAs - ok., imager ok., spectrographs didn't need to be refocued but I did refocused r1 since it was at the edge of tolerance (see the problem section). MIGs for M2 were off spec, after I did home M2 MIGs didn't change much, we decided to home M2 again, but results were the same: A,B,C axies for M2 were about 0.3 off the nominal values (see mirrors section). Galils did not reboot, and nothing has been done today with M2 - so, why today's values are different than yesturday's? Telescope ok. Night: ------ 02:01Z We pulled the enclosure. The telescope seemed slightly out of balance with the counterweights at 250 pounds, so we set them to 270. Early in the night when Steph issued a SOP command, Landru locked-up. Craig was here and rebooted it. We are not sure why Landru had this promlem. As Steph went to the first spectro field of the night, the FF fuse opened. We replaced the fuse and solved the problem. 02:49Z Cart 01, plate 1403 (chunk70). Guide stars on goToField. During the focus sweep, the guider timed out. This plate was declared done in three exposures. 04:22Z Cart 04, plate 1025 (lowz45). We had guide stars after we cleared the offsets. We lost a few minutes between this plate and the previous one because Steph had to setup a new SOP session on Landru since it crashed earlier. During the second exposure, the guider failed; the error message is in the problems section. To recover, we stopped the guider and restarted it. Four exposures were completed, however, the plate was not declared done. 06:11Z Cart 06, plate 1108 (merged48). We had guide stars on goToField. We are having some guider problems; it stopped working when we issued a scale change. During the first science exposure, the guider failed again. And ,we got the same error message as with the previous exposure. During the second science exposure, the guider stopped several times and had to be restarted. This plate was declared done after four exposures. Steph noticed that the guider seemed to have less problems after she shutdown some of the programs that were using Java animation. 08:18Z Cart 07, plate 1138 (rockosi50). We had guide stars after the offsets were cleared. We completed four exposures but were unable to declare this plate done. 10:18Z Cart 09, plate 1144 (rockosi50). Guide stars on goToField. We completed one exposure before twilight started. 11:05Z Enclosure on, doors closed, fans off. We think we could have finished a couple more plates tonight if the spectrograph was more stable, and we were not having to deal with the flop problem in r1. Mozilla was crashing on monxMike, so Craig installed Firefox on Landru. It seemed to work okay for a while, but Firefox had some prombems later in the night. Steph thinks that this problem and the guider issues may be related. endNight finished and was happy. Misc. Watcher messages in problem section. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 4763 22:49Z 23:07Z 100 O 10.23 14.74 44 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 27703, plate -9999 ------- 22:48 27703 0.0 bias ----- sequence 27704, plate 1144 ------- 22:57 27704 0.0 bias 23:03 27705 10.0 flat 23:06 27706 10.0 flat 23:09 27707 2.1 arc 23:12 27708 2.1 arc 23:31 27709 2.1 arc 23:33 27710 2.1 arc 23:43 27711 2.1 arc 23:45 27712 2.1 arc ----- sequence 27713, plate -9999 ------- 01:29 27713 0.0 bias 01:40 27714 0.0 bias 02:36 27715 0.0 bias ----- sequence 27716, plate 1403 ------- 03:12 27716 10.0 flat 03:15 27717 2.0 arc 03:34 27718 900.1 target 03:52 27719 900.1 target 04:11 27720 900.1 target ----- sequence 27721, plate -9999 ------- 04:21 27721 0.0 bias ----- sequence 27722, plate 1025 ------- 04:47 27722 10.0 flat 04:50 27723 2.0 arc 05:11 27724 900.1 target 05:29 27725 900.1 target 05:47 27726 900.1 target 06:05 27727 900.1 target ----- sequence 27728, plate 1108 ------- 06:28 27728 10.0 flat 06:31 27729 2.0 arc 06:51 27730 900.1 target 07:14 27731 1200.1 target 07:38 27732 1200.1 target 08:06 27733 1500.1 target 08:09 27734 10.0 flat 08:11 27735 2.0 arc ----- sequence 27736, plate 1138 ------- 08:29 27736 10.0 flat 08:31 27737 2.0 arc 08:51 27738 900.1 target 09:14 27739 1200.1 target 09:36 27740 1200.1 target 10:04 27741 1500.1 target 10:07 27742 10.0 flat 10:10 27743 2.0 arc ----- sequence 27744, plate 1144 ------- 10:21 27744 10.0 flat 10:24 27745 2.0 arc 10:52 27746 1200.1 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 03:18Z 1 1403 -99.09 0.0050 68.77 0.0003 94.84 -0.0206 1.000170 04:55Z 4 1025 46.60 0.0009 46.73 -0.0009 217.56 0.0090 0.999910 06:38Z 6 1108 23.21 0.0033 54.91 0.0005 199.26 0.0164 1.000290 08:34Z 7 1138 -11.59 0.0004 56.66 -0.0004 170.17 -0.0101 1.000270 10:33Z 9 1144 -53.38 0.0025 42.84 -0.0002 137.46 -0.0106 1.000290 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL 7687 Stays: JL 7688 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 03:18Z 1 1403 1.00017 -1679 -257 0 -99 68.8 13.8 20 82 gdr 1.2 04:54Z 4 1025 0.99991 889 1953 25 47 46.7 13.1 16 81 gdr 1.4 06:37Z 6 1108 1.00029 -2863 -1204 0 23 54.9 12.6 14 75 gdr 1.8 08:34Z 7 1138 1.00027 -2665 -1146 -100 -11 56.7 14.4 7 89 gdr 1.4 10:32Z 9 1144 1.00029 -2863 -1266 -100 -53 43.0 14.6 2 22 gdr 1.2 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:12Z 67 35 3 88 (E) 54646 - 22:44Z 64 34 9 81 (E) 61817 - 23:15Z 66 31 11 86 (E) 53141 - 23:48Z 62 33 12 107 (ESE) 62198 - 00:19Z 61 34 15 82 (E) 60032 - 00:53Z 61 40 11 83 (E) 59475 - 01:25Z 59 43 11 92 (E) 62616 - 01:43Z 58 44 13 80 (E) 67384 - 02:14Z 57 45 16 81 (E) 76144 - 02:44Z 57 46 19 73 (ENE) 80956 - 03:16Z 56 46 19 80 (E) 86096 - 03:47Z 56 46 18 85 (E) 83453 - 04:17Z 55 48 14 79 (E) 86385 - 04:47Z 55 47 14 82 (E) 86879 - 05:17Z 55 46 14 83 (E) 86629 - 05:48Z 54 46 12 82 (E) 87207 - 06:20Z 54 47 11 84 (E) 91164 - 06:52Z 54 47 10 87 (E) 93001 - 07:25Z 56 41 8 83 (E) 93716 - 07:57Z 58 28 6 77 (ENE) 92696 - 08:29Z 57 32 7 86 (E) 88488 - 09:01Z 58 29 7 102 (ESE) 87540 - 09:31Z 58 28 3 93 (E) 90982 - 10:02Z 59 23 1 65 (ENE) 84497 - 10:32Z 58 26 2 21 (NNE) 82079 - 11:02Z 56 28 6 11 (N) 84031 - ---=== TELESCOPE STATUS ===--- 01:15Z Doors opened, fans on, cart 01 mounted. 02:01Z Enclosure off 11:05Z Enclosure on, fans off, louvers closed Telescope at 121,30,00 Cart 09 mounted LN2 connected and turned on; spectros were force filled. CWs 270lbs Imager 182lbs 20psi Spectro 269lbs 18psi ---=== SOFTWARE USED ===--- IOP/SOP: v3_140_0 Watcher: v2_29_0 MCP: v5_25_0 TPM: tpm_v2_42_0 AstroDa: v14_47 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v1_92 SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5270 0.5260 Axial B 0.7030 0.7130 Axial C 0.9990 0.9990 Trans D -8.6820 -8.6780 Lateral E 11.5443 11.5062 Lateral F 11.5062 11.4808 GALILS Commanded: 4850. -3650. 4050. -6150. 40200. 39150. Actual: 4837. -3652. 4065. -6127. 40183. 39144. SETMIR VALUES PriDesOrient: 0.00 -3.60 26.90 1470.40 198.00 PriOrient: 0.00 -3.80 26.76 1470.93 196.30 SECONDARY: ---------- Focus: 0.00 Air Temp.: 18.9 Alt.: 30.198704 MIGS TONIGHT NOMINAL Axial A 1.4980 1.7810 Axial B 1.1080 1.3910 Axial C 1.0880 1.3710 Trans D 0.0640 0.0630 GALILS Commanded: 1628606. 1548166. 1574484. -3900. -8750. Actual: 1628641. 1548187. 1574504. -3899. -8726. SETMIR VALUES SecDesOrient: 1257.00 0.00 -25.00 0.00 146.27 SecOrient: 1256.98 0.00 -25.00 0.27 146.56 After 1-st homing: ------------------ PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5270 0.5260 Axial B 0.7030 0.7130 Axial C 0.9990 0.9990 Trans D -8.6820 -8.6780 Lateral E 11.5570 11.5062 Lateral F 11.5189 11.4808 GALILS Commanded: 4850. -3650. 4050. -6150. 40200. 39150. Actual: 4837. -3652. 4065. -6127. 40183. 39144. SETMIR VALUES PriDesOrient: 0.00 -3.60 26.90 1470.40 198.00 PriOrient: 0.00 -3.80 26.76 1470.93 196.30 SECONDARY: ---------- Focus: 0.00 Air Temp.: 16.6 Alt.: 30.346770 MIGS TONIGHT NOMINAL Axial A 1.5030 1.7810 Axial B 1.1090 1.3910 Axial C 1.0850 1.3710 Trans D 0.0710 0.0630 GALILS Commanded: 1628688. 1548014. 1574440. -3900. -8750. Actual: 1628641. 1548187. 1574504. -3899. -8726. SETMIR VALUES SecDesOrient: 1257.00 0.00 -25.00 0.00 146.27 SecOrient: 1256.95 -0.01 -25.07 0.22 146.57 After 2-nd homing: ------------------ PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5270 0.5260 Axial B 0.7020 0.7130 Axial C 0.9990 0.9990 Trans D -8.6810 -8.6780 Lateral E 11.5570 11.5062 Lateral F 11.5189 11.4808 GALILS Commanded: 4850. -3650. 4050. -6150. 40200. 39150. Actual: 4837. -3652. 4065. -6127. 40183. 39144. SETMIR VALUES PriDesOrient: 0.00 -3.60 26.90 1470.40 198.00 PriOrient: 0.00 -3.80 26.76 1470.93 196.30 SECONDARY: ---------- Focus: 0.00 Air Temp.: 15.6 Alt.: 30.194383 MIGS TONIGHT NOMINAL Axial A 1.5020 1.7810 Axial B 1.1090 1.3910 Axial C 1.0860 1.3710 Trans D 0.0670 0.0630 GALILS Commanded: 1628669. 1548235. 1574496. -3900. -8750. Actual: 1628641. 1548187. 1574504. -3899. -8726. SETMIR VALUES SecDesOrient: 1257.00 0.00 -25.00 0.00 146.27 SecOrient: 1257.02 0.02 -25.00 0.27 146.55 ---=== PROBLEMS IN DETAIL ===--- Spectrographs were in spec, however r1 was at the spec border, so I piston collimator (checks done at 14.5 C). Before collimation: sp1> mechSend s spMechVersion v2_5_0 Bootup 719771169 SpectroID 1 SlitID 41 Air On Shutter_open_sensor Off Shutter_closed_sensor On Left_open_sensor Off Left_closed_sensor On Right_open_sensor Off Right_closed_sensor On Coll_motor_A 2824 Coll_motor_B 3804 Coll_motor_C 1924 Requested_exp.time 0.0 Exp_time_left 0.0 Last_exp.time 2.1 Exp_state None Shutter_open_transit .70 Shutter_close_transit .52 Coll_motor_A_status 0x84 Coll_motor_B_status 0x84 Coll_motor_C_status 0x85 sp1> mechPiston -2400 after collimation: ------------------ sp1> mechSend s spMechVersion v2_5_0 Bootup 719771169 SpectroID 1 SlitID 41 Air On Shutter_open_sensor Off Shutter_closed_sensor On Left_open_sensor Off Left_closed_sensor On Right_open_sensor Off Right_closed_sensor On Coll_motor_A 425 Coll_motor_B 1398 Coll_motor_C -476 Requested_exp.time 0.0 Exp_time_left 0.0 Last_exp.time 2.1 Exp_state None Shutter_open_transit .70 Shutter_close_transit .52 Coll_motor_A_status 0x85 Coll_motor_B_status 0x84 Coll_motor_C_status 0x85 sp1> p2> mechSend s spMechVersion v2_5_0 Bootup 719771169 SpectroID 2 SlitID 9 Air On Shutter_open_sensor Off Shutter_closed_sensor On Left_open_sensor Off Left_closed_sensor On Right_open_sensor Off Right_closed_sensor On Coll_motor_A 2259 Coll_motor_B 52 Coll_motor_C -641 Requested_exp.time 0.0 Exp_time_left 0.0 Last_exp.time 2.1 Exp_state None Shutter_open_transit .72 Shutter_close_transit .62 Coll_motor_A_status 0x85 Coll_motor_B_status 0x84 Coll_motor_C_status 0x85 sp2> ---------------------------------------------- --== Misc Watcher Messages ==-- Text="failed to move one or more axes":AxisNoTrack:"TTT" (2004-08-10 02:18:47Z) Text="Could not queue a write to user:message output dies": (2004-08-10 02:31:07Z) Facil="SYSTEM":VMSMsg:"DEVOFFLINE" (2004-08-10 02:31:07Z) Facil="SYSTEM":VMSMsg:"HANGUP" (2004-08-10 02:31:07Z) Modu="axe i SchMove" (2004-08-10 02:19:14Z) Modu="axe o Init" (2004-08-10 02:19:14Z) Modu="exe Track" (2004-08-10 02:19:14Z) Text="Write to user failed:output terminates to user8": (2004-08-10 02:31:07Z) Modu="prt ReadMatch" (2004-08-10 02:41:04Z) Modu="prt ReadReplySet" (2004-08-10 02:18:47Z) Modu="prt Init Forth" (2004-08-10 02:41:04Z) dewar (Imager Dewar Scale Communications Fault:: 2004-08-10 01:05:58Z) tracking (Excessive Az Tracking Error:: 2004-08-10 02:01:29Z) m2 mig (Communications Fault: 2004-08-10 00:46:35Z) azimuth (correction for azimuth 596356 is too large (2004-08-10 06:11:07Z sdssmcp): 2004-08-10 06:11:09Z ) --------------------------------------------------------- --== Guider Problem ==-- We got the following error message when the SOP guider failed: expected floating-point number but got "" expected floating-point number but got "" while executing "format " %12.6f" $guiderData($expNum,AltMount) " invoked from within "puts $outFile [format " %12.6f" $guiderData($expNum,AltMount) ]..." (procedure "writeGuideIter" line 69) invoked from within "writeGuideIter" (procedure "guiderLoop" line 137) invoked from within "guiderLoop $verbose $boresight" invoked from within "if {[checkForNewImage 1 0] != ""} { schedule checkAndGuide {} 0 guiderLoop $verbose $boresight set guiderInfo(waiting) 1 schedule asyncCheck [list ..." (procedure "checkAndGuide" line 14) invoked from within "checkAndGuide 0 0 15" ("eval" body line 1) invoked from within "eval $procname $args" checkAndGuide is scheduled, further errors sent to murmur